UT Nov 17 2021 - E/D Night




Night Info:

  • Staring Config - LBTI
  • Sunset 17:23
  • Twilight Ends 18:43
  • Moon sets 05:20
  • Twilight Starts 05:26
  • Sunrise 06:47

Night Plan:

  • LBTI Observations of Pallas - target sets at 22hr local time
    • Program can withstand some clouds and seeing up to 1.2"-1.5"
  • Reconfigure to MODS-Bino 22hr-22:15hr
  • Science M31 Nova 22:15hr-23:30hr
    • Conditions - some clouds and seeing up to 1.5"
  • Reconfigure to LBCs 23:30hr-00hr
  • TMS Testing 00hr-03hr
  • Reconfigure LUCI 03hr-03:30hr
  • LUCI Sub-frame tests / Poor Weather Observing

Staff

  • OSA: J. Williams
  • ISp: E. Solheid
  • Mtn Manager: M.Wagner
  • ISA; B. Rothberg, R.T.Gatto

Night Log

Closed Dome

  • UT 23:09 (-1 day) - waking up MODS, test images
    • all ok
  • UT 23:28 (-1 day) - LUCI GUIs started up, init all
    • all ok
  • UT 23:31 (-1 day) - powering up LBCs
    • UT 23:35 (-1 day) test biases - all OK
  • UT 23:33 (-1 day) - sky is completely overcast
  • UT 00:12 - sky still completely overcast
  • UT 00:18 - Opening enclosure
UT 00:40 - Lovely sunset

webcam.jpg

LBTI - Pallas Observations

  • Trying a pointing star first (V=2.5mag)
  • Can see Pallas in viewer, trying Bin2 (not bright enough for Bin 1)
  • can't close the loop. most of the pupils is just sky background
  • UT 01:05 - will wait and contnue to try as sky darkens and/or clouds thin out
  • Closed loop on Pallas - need to apply GOPT and NCPA
  • UT 01:15 - starting 10min of data collection
  • UT 01:27 - taking science data
  • UT 02:11 - moving to reference star
  • UT 02:44 - Amali reports something not right with GOpt
  • UT 02:55 - moving to brighter star to diagnose issues
  • UT 03:18 - first PSF star had a companion, was not a problem with the WFS - new star found, data being taken
  • UT 03:24 - stopping now - clouds thickening up

MODS - M31 Nova

  • UT 03:25 - reconfiguring for bino-MODS
  • First target - Feige 110 - SpecStd
  • UT 03:39 - pointing check 3rd mag star
  • UT 03:42 - pointing check 6th mag star
  • UT 03:45 - collimation check
  • guiders 1.1"
  • UT 03:37 - preset to Feige 110
    • guiders 1.05"
    • M2 offsetxy 4.393 9.349 rel
    • M1 offsetxy -0.909 12.124 rel
    • getting too many DS9 windwos open error from modsAlign
      • killed all DS9s on OBS2, restarted modsDisps then modsAlign, was happy
    • M2 tweak offsetxy -0.359 -0.083 rel
    • M1 tweak offsetxy 0.149 0.222 rel
  • UT 03:56 - starting spectra
  • UT 03:59 - looks like the thicker clouds are heading out to the east
    • seeing 1.5" on guiders
  • given the seeing and thin clouds - will use a 1.2" slit
  • UT 04:09 - sending preset to M31 Nova PA=0, 1.2" slit
    • seeing at target ~1"
    • back up to 1.2"-1.3"
    • M1 offsetxy -0.276 10.447 rel
    • M2 offsetxy 4.100 8.338 rel
    • both thru-slits look OK
  • UT 04:22 - starting spectra
    • UT 04:39 - 0.9"-1"
    • object significantly fainter, can still see bright h-alpha in Red channel
    • UT 04:54 - guiders show 0.7"-0.8" now, but does pop back up to 1"
    • UT 05:29 - will take one additional 1200sec exposure

MODS WFS CutOut Tests

  • Run through each quadrant WFS script
  • UT 05:52 - sending lower left script
    • M2R IMCS timed out - hit retry
    • dither that triggered it was +/-1.5" DET Y
    • M1R shows a jump in image, M1B is OK
    • series of dots in M1R
    • otherwise no WFS cutout issue
    • did another 0, 1.5" DETXY offset, WFS cutout did not appear, took exposures on both MODS
    • no jumps
    • looks like debris in MODS1, both channels
  • UT 06:12 - sending lower right with offsets
    • no issues
    • UT 06:22 - still doing exposure done cleanup, blue expdone fixed it
  • UT 06:24 - sending upper left with offsets
    • for all tests it does look like GCS says WFS Paused (green font) while the exposure starts
    • no issues for this quadrant
  • UT 06:35 - sending upper right with offsets
    • WFS cutouts look OK on both sides
    • no change in cutout with dithers
  • UT 06:45 - Reconfiguring for LBCs

LBC/TMS Testing

Summary

TMS was run in active mode last night from about 07:30 until 12:45 UT. An old reference from 20211115 UT was used to help the initial collimation, and it did, as the first pupils only had Z4 outside of the convergence criteria. The ambient temperature last night around 07:30 was ~1.8 and it dropped by almost 1 deg, to ~0.9 deg C, over the course of the LBC/TMS testing. The temperature was cooler than the night of 20211115 UT (~6 C) when the previous reference had been taken.

We alternated between observing fields at high (~75-80 deg) and low (~35 deg) elevation and bracketed the observations by a pair of extra-focal pupil images. FPIA was run at the start, ~07:45, and not again until the end of the night, ~12:45. The image quality looked good throughout these 5 hours and at the end of the night, the first and only pair of pupil images were within the convergence criteria for the estimate seeing (1.09" blue and 0.85" red). The cumulative thermal Z4 corrections were almost -1000 nm.

There were at least two LBCR images which showed image jumps: lbcr.20211117.115008.fits and lbcr.20211117.121603.fits; and a few red images showed elongation. There was a period when the LBCB image quality looked as good as or better than the LBCR IQ. The seeing was variable, going from 2" at one low elevation observation to ~0.8". All in all, it was a good half-night of testing.

One noteworthy issue is the pointing and copointing. Again, we found that the two sides were not well pointed or copointed and two times after commanding the pointing model changes (dIE and dCA) and the optics offsets (X,Y,RX, RY), the preset was lost. Nevertheless the updates were made and we carried on observing the fields.

The procedure pointed out something I had not noticed before - the instrument focus offset is not updated until after the slew has completed. Usually we are not slewing ot a new field when M1 is defocussed by 0.8mm. Loading a copointing OB and then immediately stopping it did not give time for the mirror to move to focus. When the mirror is 0.8mm below focus, it seems like, although TMS does not send a Z correction, it does adjust the mirror X,Y, RX, RY positions - is that correct? To be sure, we watched for updates to the global offset before taking the extra-focal pupil image at the new position.

Log

  • Copied the reference from Nov 15 UT
[telescope@robs 20211115]$ cp tms_sx_ref.dat /tmp
[telescope@robs 20211115]$ cp tms_dx_ref.dat /tmp
[telescope@robs 20211115]$ cp telescope_left_tmsoffset.dat /tmp
[telescope@robs 20211115]$ cp telescope_right_tmsoffset.dat /tmp

07:35(about) Starting tms_loop.pl

07:40 Slewing to copointing field for 2013 MW6 (elevation 85 deg and setting).

07:46 dofpia, after waiting for TMS to settle First corrections are pretty well collimated except for z4
  5592   -124     54   -175    293    146      4   1.36   2.21     red
  3851   -211   -370     92     59   -163    -41   1.38   1.53    blue

07:52 set reference (./lbc_set_tms_offset.py -s 2) and start tms_loop.py.
  • Note: the pupil images were not well copointed

07:54 Copointing
COPOINTING: B=75439 R=75434
Pointing updates: delta_IE = 34.50", delta_CA = -11.42"
  Mirror updates:   dX(mm)   dY(mm)   dRX(")   dRY(")
              SX:    0.41     0.82    17.69    -8.72
              DX:   -0.47    -0.90   -19.34     9.98

Preset was lost on both sides.

08:02 Took the pair of copointing images again and ran lbcrangebal again to verify that the corrections were sent despite the lost preset. They were.

08:04 Starting observations on 2013 MW6 (elevation 79 deg).
  • face-back temperature differences are -0.061 (SX) and -0.156 (DX). The glass is about 0.3 (SX) and 0.6 (DX) warmer than ambient.
  • FWHM on first pair of images: 7.5 pix (1.7", Blue) and 6.8 (1.5", Red)
  • 60s exposures - backgrounds are pretty high, but OK ~15k blue and 12k red
  • DIMM seeing 2"
  • 08:37 FWHM 6.5 (Blue) and 5.5 (Red)
  • Script ended at 08:43

08:43 Control-C TMS loop and take extra-focal pupil image.
  • ./addsph.py
  • RB_fastextra4 : b 084619 and r 084615
  • ./remsph.py

08:51 Slew to low elevation field (2021 UQ1)
  • Wait for group rotation to change (definitely saw it change on DX and think so on SX)
  • ./addsph.py
  • RB_fastextra4 : b 085805 r 085800 (use these and not the 2nd pair which were taken after z4/z11 removed)
  • ./remsph.py

09:00 Taking copointing images (OB didn't load the first time)
COPOINTING: B=90333 R=90329
Pointing updates: delta_IE = -9.23", delta_CA = 14.22"
  Mirror updates:   dX(mm)   dY(mm)   dRX(")   dRY(")
              SX:    0.07    -0.61   -13.04    -1.40
              DX:   -0.12    -1.85   -39.72     2.66
  • Copointing images are a bit elongated (blue) and round but large (red)

09:09 Started science OB for 2021 UQ1
  • Images have FWHM ~ 7 pix
  • Backgrounds 20k, 15k on b, r

09:16 Slewing to 2004 QB copointing field. Waiting until group rotation updates to pause TMS loop.
  • ./addsph.py
  • RB_fastextra4 : b 092150 and r 092145
  • ./remsph.py

09:23 Taking copointing images 092503 and
COPOINTING: B=92503 R=92456
Pointing updates: delta_IE = -1.72", delta_CA = -13.90"
  Mirror updates:   dX(mm)   dY(mm)   dRX(")   dRY(")
              SX:   -0.01     0.57    12.29     0.11
              DX:    0.08     1.41    30.23    -1.79

09:28 Restarting TMS loop

09:29 Starting OB for 2004 QB1 (object is fainter than the other two but 60 deg from moon).
  • elevation 69 deg
  • backgrounds in 60s are 15k (b) and 10k (r)
  • FWHM on first images 5.2 (1.2", b) and 4.8 (1.1", r). DIMM seeing 1.35" (zen)
  • IQ is looking a bit better on the blue than red.
  • Stopping OB after 16 images per side (16-min open-shutter)

10:01 Pausing TMS
  • addsph.py
  • Running RB_fastextra4 : b 100229 and r 100228
  • remsph.py

10:05 Slewing to GD279 copointing field (elev ~35 deg) and resuming TMS loop, waiting until group rotation changes.

10:09 Taking pair of pupil images
  • addsph.py
  • RB_fastextra4 : b 101205 101201
  • remsph.py

10:14 Taking pair of copointing images

COPOINTING: B=101503 R=101459
Pointing updates: delta_IE = -19.85", delta_CA = 18.58"
  Mirror updates:   dX(mm)   dY(mm)   dRX(")   dRY(")
              SX:    0.04     0.39     8.47    -0.82
              DX:   -0.02    -0.35    -7.56     0.52

10:21 Observing GD279 (did not start TMS loop until (10:31)
  • It seems clear from the all-sky
  • elevation 33:41
  • guide star FWHM ~2.2-2.5"
  • B & R, first images 60-sec but backgrounds are 12k (B) and 20k (R)
  • ~10:30 FWHM 9 pix (blue) and 6-7 pix (red) but we noted that TMS loop was not running
  • 10:32 DIMM 1.5" (zen)
  • Paused and resumed with 30-sec exposure times.
  • FWHM 9.5 (blue) and 7 (red)
  • 10:41 Stopped

10:42 Take pair of pupil images
  • addph.py
  • RB_fastextra4 : b 104242 r 104238
  • remsph.py
  • started TMS loop

10:44 Slewing to copointing field for sdssY (79 deg) and wait for group rotation updates.

10:48 Taking pair of pupil images
  • addsph.py
  • RB_fastextra4 : b 105102 and r 105057
  • remsph.py
  • restarting TMS loop
10:55 Taking copointing image pair * To let a tms correction happen, I stopped and restarted the copointing OB, but the correction was very small (not necessary since it was already in position - the M1 only had to move 0.8mm)
Radial star offsets from rotator centers: BLUE 35.0" and RED 50.1"

lbcrangebal:
-------------------------------------------------------------------------
COPOINTING: B=105620 R=105615
Pointing updates: delta_IE = 37.40", delta_CA = -19.97"
  Mirror updates:   dX(mm)   dY(mm)   dRX(")   dRY(")
              SX:   -0.06    -0.45    -9.71     1.36
              DX:    0.08     0.43     9.24    -1.63

11:00 Starting the observation g+r on sdssY7 (elev 81 deg)
  • It seems clear from the all-sky
  • 60-sec g+r (UUs + izY)
  • 80 deg from moon
  • FWHM ~4pix blue and 3pix red, later (11:18) 4.5 b and 3.8 r. DIMM has been reading 1.1"-1.28" (zen)
  • 11:25 Stopping the OB

11:26 Pausing TMS to take pupil images
  • addsph
  • RB_fastextra4 : b 112814 and r 112753
  • remsph
  • Start TMS loop

11:30 Slewing to 2013 MW6 copointing field (elev 36 deg)
  • I've been noticing that M1 doesn't move (PSF GUI instrument offset remains at -0.8) until the slew is complete.-0.8mm
  • Waiting for group rotation to update

11:34 Taking pupil images (after update to group rotation)
  • Paused TMS
  • addsph.py
  • RB_fastextra4 : b 113530 r 113525
  • remsph.py

11:38 Taking pair of copointing images
  • restarted TMS loop
Radial star offsets from rotator centers: BLUE 32.7" and RED 48.5"

lbcrangebal:
-------------------------------------------------------------------------
COPOINTING: B=113932 R=113928
Pointing updates: delta_IE = -38.29", delta_CA = 13.33"
  Mirror updates:   dX(mm)   dY(mm)   dRX(")   dRY(")
              SX:    0.07     0.43     9.16    -1.50
              DX:   -0.07    -0.41    -8.72     1.45

11:43 Observing 2013 MW6 (elev 34 deg)
  • The red image in the first pair (114346) looked a bit elongated, but the 2nd red image was better
  • FHWM ~ 4.7pix (blue) and 3.7pix (0.83", red), DIMM ~ 0.73" (zen)
  • Red 115008 shows some trails
  • 11:58 Stopped
11:59 Taking pair of pupil images
  • Stop TMS loop
  • addsph
  • RB_fastextra4 : b 115939 and r 115933
  • remsph
  • Resume loop

12:08 Slewed to field 2004 QB (elev 52). Waited for TMS to update group rotation.

12:09 Take pair of pupil images
  • addsph.py
  • RB_fastextra4
  • remsph.py
  • Starting loop

12:12 Taking pair of copointing images

Radial star offsets from rotator centers: BLUE 5.6" and RED 11.3"

lbcrangebal:
COPOINTING: B=121319 R=121316 Pointing updates: delta_IE = 3.63", delta_CA = 6.91" Mirror updates: dX(mm) dY(mm) dRX(") dRY(") SX: -0.06 -0.16 -3.37 1.35 DX: 0.06 0.19 4.06 -1.36

12:15 Starting 2004 QB observation (elev 50 deg)
  • First red image 121603 is trailed/jump but first blue image 121610 looks good (FWHM 4.1 pix)
  • 121808 and 121804 have 4 pix (b) and 3.8 (r) and look round.
  • 12:41 Stopping the OB

12:41 Taking a pair of pupil images here (elev now 44:50)
  • Pause TMS
  • addsph
  • RB_fastextra4 : b
  • remsph
  • resuming TMS

12:44 Slewing to sdssY8 copoint (elev 80)
  • Waiting for TMS corrections
  • Running dofpia
  • First pupils were converged
-328   -116   -297    -37    122   -104    -17   1.09   0.20    blue
-136    -36     51    166   -117    -45    -29   0.85   0.14     red
  • Thermal corrections are Z4 = -965 nm (SX) and -973 nm(DX)

13:01 Observing sdssY8 (in twilight! and with some cirrus...)
  • LBCR has vertical streaks on chips 1 and 3 - readout artifacts
    • lbcr.20211117.130309.fits and lbcr.20211117.130402.fits
      • These were the first images after the copointing field was observed. The copointing OB used only chip 2.
    • Turning off/on systems (including controller) resolved the problem

13:15 Slewing to the field for twilight sky flats

13:18 Starting g + r flats
  • b 132112 is "scrambled"

13:22 Starting B + R flats --- not so good. Hitting pause did not "take" when I wanted and upon resuming after a "stop", the rotator had a long way to go on B but not R?

13:30 Trying S + Y flats
  • b 133203 is "scrambled"

13:34 Josh is closing up

RB_Science crashed:

---- receiving 'lbcr.20211117.083155.fits' [frame 2] (5.0s/95.0s)
      object:2013 MW6 17No | filter:r-SLOAN | exptime:30.24 | pa_pnt:-0.09 | dithoffx:4 | dithoffy:1

Unknown error! Please restart (and report the following messages).
------------------------------------------------------------
Traceback (most recent call last):
  File "/lbt/lbto/supportscripts/ObserverSupport/RB_Science.py", line 951, in display
    sched.run()
  File "/lbt/lbto/supportscripts/ObserverSupport/RB_Science.py", line 611, in run
    t.run()
  File "/lbt/lbto/supportscripts/ObserverSupport/RB_Science.py", line 860, in run
    [f'frame {self.__frameno}']+self.__cmds.expand(img))
  File "/lbt/lbto/supportscripts/ObserverSupport/RB_Science.py", line 353, in set
    raise err # temporarily rethrow the exception for now.
  File "/lbt/lbto/supportscripts/ObserverSupport/RB_Science.py", line 351, in set
    self.__ds9.set(cmd)
  File "/usr/lib64/python3.7/site-packages/pyds9/pyds9.py", line 646, in set
    self._selftest()
  File "/usr/lib64/python3.7/site-packages/pyds9/pyds9.py", line 571, in _selftest
    1):
  File "/usr/lib64/python3.7/site-packages/pyds9/xpa.py", line 209, in xpaaccess
    raise ValueError(errmsg)
ValueError: XPA$ERROR: no response from server during handshake (?:?)

-- BarryRothberg - 16 Nov 2021
Topic revision: r9 - 17 Nov 2021, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback